OpenStudy (anonymous):

Find the x-intercepts for the parabola defined by the equation below. y = 3x2 + 18x + 15 (-1, 0) and (-5, 0) (5, 0) and (6, 0) (0, -1) and (0, -5) (0, 5) and (0, 6)

OpenStudy (muzzack):

No Solutions found Rearrange : Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : y-(3*x^2+18*x+15)=0 Step by step solution : Step 1 : Simplify y - 3x2+18x+15 Checking for a perfect cube : 1.1 y-3x2-18x-15 is not a perfect cube No factorization could be found Equation at the end of step 1 : y - 3x2 - 18x - 15 = 0 Step 2 : Solve y-3x2-18x-15 = 0 Solving a Single Variable Equation : 2.1 Solve y-3x2-18x-15 = 0 In this type of equations, having more than one variable (unknown), you have to specify for which variable you want the equation solved. We shall not handle this type of equations at this time. No Solutions found

OpenStudy (anonymous):

(-1, 0) and (-5, 0) (5, 0) and (6, 0) (0, -1) and (0, -5) (0, 5) and (0, 6) these are the choices @Muzzack

OpenStudy (muzzack):

hmm im sorry

OpenStudy (imstuck):

This is an easy one! When you are solving for the x intercepts, you will factor the polynomial they gave you and find where x = 0. When you sove this, you can factor out a 3 to leave y = 3(x^2 + 6x 5). Factor that and you get y = 3(x + 1)(x + 5). Setting each of these factors equal to 0 you get x = -1 and x = -5. That is where your polynomial goes through the x axis (aka, the x intercepts of your graph). Get it?

OpenStudy (vishweshshrimali5):

x-intercepts mean the points where the graph of a function intersect the x-axis. To find out these points put y = 0.

OpenStudy (imstuck):

Sorry that one polynomial left out a "+" sign. It's 3(x^2 + 6x + 5). But the rest of it is the same and is correct.